# Generating Molecular Conformer Fields

### v2

@@ 6 Conclusions
In this paper we introduced MCF, where we formulate the problem of molecular conformer generation as learning a distribution over conformation fields on molecular graphs. MCF is formulated specifically as a diffusion generative model over fields, which achieves state-of-the-art performance compared with baselines with strong inductive bias when evaluated on molecular generation across different benchmarks. Notably, MCF achieves these results without explicitly modeling geometric properties of molecules like torsional angles, which makes it simpler to understand and scale. MCF presents an exciting avenue for future research on scaling conformer generation to proteins and other macro molecular structures. We further discuss limitations and future works in Appendix A.1.
@@ A.1 Limitations and Future Work
While MCF shows competitive performance in molecular conformer generation, it does encounter limitations and potential improvements for future explorations. One limitation is that our proposed method is computationally expensive. Extensive computations first stem from the Transformer-based (Vaswani et al., 2017) score network. In MCF, we use a PerceiverIO (Jaegle et al., 2022) as score network, an efficient transformer that allows for sub-quadratic compute. Other efficient transformer architectures and tricks like (Dao et al., 2022) can be used to improve training efficiency. The other factor is computational cost during inference. In MCF, we iterate 1000 timesteps to sample a conformer following DDPM (Ho et al., 2020). Experiments in Appendix A.4 further show that efficient sampling strategies, i.e.  DDIM (Song et al., 2021), can help significantly increase inference efficiency while maintain high-quality in sampled conformers. Other efficient variants of diffusion models like consistency model (Song et al., 2023) as well as distillation approaches (Berthelot et al., 2023) may be adapted to further decrease the sampling to single step. Also, recent works have demonstrated that diffusion generative model can generate samples following Boltzmann distributions when provided with Boltzmann-distributed training data (Arts et al., 2023). Driven by this, our proposed MCF can be adapted to generate molecular conformers that follow Boltzmann distributions when trained with corresponding data. Besides, recent flow matching generative model (Lipman et al., 2022) provides the flexibility of mapping between arbitrary distributions and access to exact log-likelihood estimation. Integrating flow matching framework could help sample molecular conformer from Boltzmann distribution instead of standard Gaussian.
Another limitation could be the fact that MCF does not explicitly model the equivariance of the molecular system. Though experiments in the paper have demonstrated that equivariance may not be necessary to achieve competitive performance on conformer generation. MCF may not perform as well as conformer generation when applied to problems with limited data or related to sequential problems like molecular dynamics (MD) simulations. In future work, we plan to extend MCF to conditional inference. For example, molecular docking can be formulated as conformer generation problem conditioned on proteins (Corso et al., 2022). Also, current framework can be expanded to de novo drug designs where no molecule information is provided (Hoogeboom et al., 2022). Besides, scaling up our model to large molecules, like proteins, can be of great interest. MCF by nature provides the flexibility to generate from partially observed sample, which can be suitable for designing proteins with known functional motifs (Watson et al., 2023).
@@ A.2 Implementation details
In this section we describe implementation details for all our experiments. We also provide hyper-parameters and settings for the implementation of the score field network $\epsilon_{\theta}$ and compute used for each experiment in the paper.
@@ A.2.1 Score Field Network implementation details
The time-step $t$ is incorporated into the score computation by concatenating a positional embedding representation of $t$ to the context and query sets. The specific PerceiverIO settings used in all quantitatively evaluated experiments are presented in Tab. 5. Practically, the MCF network consists of 6 transformer blocks, each containing 1 cross-attention layer and 2 self-attention layers. Each of these layers has 4 attention heads. An Adam (Kingma & Ba, 2015) optimizer is employed during training with a learning rate of $41e-$ 4. We use EMA with a decay of $0.9999$. A modified version of the publicly available repository is used for PerceiverIO  [^1] https://huggingface.co/docs/transformers/model_doc/perceiver. Since molecules have different number of atoms, we set the number of context and query sets as the number of atoms during training and inference.
Hyper-parameter GEOM-QM9 GEOM-DRUGS Ablation GEOM-QM9
#eigenfuncs ($k$) $28$ $32$ $28$
#freq pos. embed $t$ $64$ $64$ $64$
#latents $512$ $512$ $512$
#dim latents $512$ $512$ $256$
#model dim $768$ $1024$ $768$
#blocks $6$ $6$ $6$
#dec blocks $2$ $2$ $2$
#self attends per block $2$ $2$ $2$
#self attention heads $4$ $4$ $4$
#cross attention heads $4$ $4$ $4$
batch size $64$ $128$ $64$
lr $41e-$ 4 $41e-$ 4 $41e-$ 4
epochs $250$ $300$ $250$
@@ A.2.3 Compute
Each model was trained on an machine with 8 Nvidia A100 GPUs, we trained models for 500 epochs.
@@ A.3.1 Ablation experiments
In this section we provide an ablation study over the key design choices of MCF. We run all our ablation experiments on the GEOM-QM9 dataset following the settings in GeoMol (Ganea et al., 2021) and Torsional Diffusion (Jing et al., 2022) and described in Sect. 5.2. In particular we study: (i) how does performance behave as a function of the number of Laplacian eigenvectors used in $\varphi(v)$. (ii) How does the model perform without atom features (e.g. how predictable conformers are given only the graph topology, without using atom features). Results in Tab. 7 show that the graph topology $\mathcal{G}$ encodes a surprising amount of information for sampling reasonable conformers in GEOM-QM9, as shown in row 2. In addition, we show how performance of MCF changes as a function of the number of eigen-functions $k$. Interestingly, with as few as $k=2$ eigen-functions MCF is able to generate consistent accurate conformer.
Recall Precision
Coverage $\uparrow$ AMR $\downarrow$ Coverage $\uparrow$ AMR $\downarrow$
$k$ Atom Features mean median mean median mean median mean median
28 YES 94.86 100.00 0.125 0.081 91.49 100.00 0.175 0.122
28 NO 90.70 100.00 0.187 0.124 79.82 93.86 0.295 0.213
16 YES 94.87 100.00 0.139 0.093 87.54 100.00 0.220 0.151
8 YES 94.28 100.00 0.162 0.109 84.27 100.00 0.261 0.208
4 YES 94.57 100.00 0.145 0.093 86.83 100.00 0.225 0.151
2 YES 93.15 100.00 0.152 0.088 86.97 100.00 0.211 0.138
Table 7: Experiments on GEOM-QM9 with different numbers of eigenvectors. In these experiments, we use a smaller model than what we have in Table 1
@@ A.3.2 Architectural choices
To further investigate the design choices of architecture in proposed MCF, we include additional experiments on GEOM-QM9 as shown in Tab. 8. To investigate the effectiveness of using Laplacian eigenvectors as positional embedding, we leverage SignNet (Lim et al., 2022) as the positional embedding, which explicitly models symmetries in eigenvectors. Using SignNet does not benefit the performance when compared with the standard MCF. Though adding edge attributes in SignNet achieves better performance than SignNet alone, the performance is still not rival. Also, it’s worth mentioning that SignNet includes graph neural networks (Xu et al., 2018) and Set Transformer (Lee et al., 2019) which makes training less efficient.
In addition, we also report results using a vanilla Transformer encoder-decoder (Vaswani et al., 2017) as the backbone instead of PerceiverIO (Jaegle et al., 2022). The Transformer contains 6 encoder layers and 6 decoder layers with 4 attention heads. Other model hyperparameters follow the same setting as listed in Tab. 5. It is indicated that Transformer is performing worse than PerceiverIO on conformer generation, which validates the design choice of architecture in MCF.
@@
Coverage $\uparrow$ AMR $\downarrow$ Coverage $\uparrow$ AMR $\downarrow$
mean median mean median mean median mean median
MCF SignNet 94.1 100.0 0.153 0.098 87.5 100.0 0.222 0.152
MCF SignNet (edge attr.) 95.3 100.0 0.143 0.091 90.2 100.0 0.197 0.135
MCF Transformer 94.3 100.0 0.159 0.111 90.7 100.0 0.202 0.136
MCF PerceiverIO 96.57 100.00 0.107 0.072 94.60 100.00 0.130 0.084
Table 8: Molecule conformer generation results with different network architectures on GEOM-QM9.
@@ A.3.3 Ensemble properties
Torsional Diff. (our eval) 3.07$\pm$2.32 0.61$\pm$0.38 1.71$\pm$1.69 4.11$\pm$7.91
MCF (ours) 1.00$\pm$0.70 0.44$\pm$0.36 1.32$\pm$1.40 1.16$\pm$2.02
@@
Figure 6: Examples of conformers with lowest and highest energies in ground truth, MCF samples, and Torsional Diffusion samples for different molecules.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Figure 7: Distributions of energies ($E$), dipole moments ($\mu$), and gap ($\delta\epsilon$) of molecules a, b, c randomly picked from GEOM-DRUGS.
@@
To further evaluate the performance on ensemble properties, we randomly pick 10 molecules from test set of GEOM-DRUGS and compare MCF with our replica Torsional Diffusion on the subset as shown in the last two rows of Tab. 9. We use the checkpoints from the public GitHub repository [^2] https://github.com/gcorso/torsional-diffusion of Torsional Diffusion to sample conformers. Unlike previous setting which only sample 32 conformers, we sample $2K$ conformers for a molecule with $K$ ground truth conformers. We report the average and standard deviation of errors over the 10 molecules. It is indicated that MCF generates samples with ensemble properties that are closer to the ground truth. Fig. 6 shows the conformers with lowest and highest energy in ground truth, MCF samples, and Torsional Diffusion samples. Fig. 7 shows the ensemble property distributions of conformers from ground truth, Torsional Diffusion, our MCF. MCF generates samples conformers with ground states close to the ground truth. Sometimes it can even find conformers with lower energy (e.g. Fig. 6(c)). However, both MCF and Torsional Diffusion can generate conformers with higher energies than ground truth.
A.4 Efficient Sampling
To reduce the computational cost of MCF in inference, we apply efficient sampling technique, i.e.  DDIM (Song et al., 2021), with significantly smaller number of sampling steps. Tab. 10 show the comparison of MCF and Torsional Diffusion (Jing et al., 2022) with different sampling steps. With 20 or 50 sampling steps, MCF achieves comparable performance as the DDPM with 1000 sampling steps and greatly surpasses Torsional Diffusion with same sampling steps. Indeed, with very small sampling steps (e.g. 3 or 5 steps), Torsional Diffusion demonstrate better performance than MCF. We attribute this to the fact that Torsional Diffusion uses a pre-processing step in which a cheminformatics method predicts local substructure in 3D spaces before applying their model. So even before inference the cheimformatics prediction provides a strong prior. It should also be noted that with small sample steps, Torsional Diffusion also suffer from low sample quality compared with the counterparts with more steps. Fig. 8 show some examples of sampled conformers from MCF with different sampling steps. Even with small sampling steps, MCF can still generate plausible 3D structures, especially for the heavy atoms.
Precision Recall
Coverage $\uparrow$ AMR $\downarrow$ Coverage $\uparrow$ AMR $\downarrow$
steps mean median mean median mean median mean median
Torsional Diff. 3 42.9 33.8 0.820 0.821 24.1 11.1 1.116 1.100
Torsional Diff. 5 58.9 63.6 0.698 0.685 35.8 26.6 0.979 0.963
Torsional Diff. 10 70.6 78.8 0.600 0.580 50.2 48.3 0.827 0.791
Torsional Diff. 20 72.7 80.0 0.582 0.565 55.2 56.9 0.778 0.729
Torsional Diff. 50 73.1 80.4 0.578 0.557 57.6 60.7 0.753 0.699
MCFDDIM 3 15.05 0.00 1.032 1.041 5.71 0.00 1.521 1.525
MCFDDIM 5 42.86 35.50 0.813 0.814 20.07 11.54 1.149 1.149
MCFDDIM 10 74.14 83.25 0.610 0.601 48.95 46.35 0.841 0.813
MCFDDIM 20 80.87 88.89 0.522 0.504 59.72 60.23 0.745 0.708
MCFDDIM 50 81.87 88.89 0.481 0.459 62.01 62.53 0.708 0.660
MCFDDIM 100 81.97 89.10 0.466 0.441 62.81 63.64 0.693 0.641
MCFDDPM 1000 81.62 89.22 0.468 0.438 61.63 62.50 0.705 0.650
Table 10: Molecule conformer generation results with different sampling strategies and timesteps on GEOM-DRUGS.

### v3

@@ 6 Conclusions
In this paper we introduced MCF, where we formulate the problem of molecular conformer generation as learning a diffusion model over functions on molecular graphs. MCF achieves state-of-the-art performance across different molecular generation benchmarks, surpassing models with hard-coded inductive biases by a large margin. Notably, MCF uses general-purpose Transformer-based score network rather than a model designed with specific inductive biases for molecules. MCF achieves superior results without explicitly modeling geometric properties of molecules like torsional angles, which makes it simpler to understand and scale. We believe MCF represents an exciting first step for future research on scaling conformer generation to proteins and other macro molecular structures. We hope our work serves as a reminder to the community to carefully consider the interplay between baking inductive biases in architectures while also considering the benefits of efficient and scalable approaches.
Impact Statement
This paper introduces a novel diffusion model based approach to generate molecular conformers, contributing to the advancement of computational chemistry and molecular modeling. The impact of this work extends to various scientific domains, enabling more accurate predictions of molecular structures and behaviors. The potential societal implications lie in the acceleration of drug discovery, materials science, and environmental studies, offering efficient solutions to complex molecular design challenges.
@@ A.1 Limitations and Future Work
While MCF shows competitive performance in molecular conformer generation, it does encounter limitations and potential improvements for future explorations. One limitation is that our proposed method is computationally expensive. Extensive computations first stem from the Transformer-based (Vaswani et al., 2017) score network. In MCF, we use a PerceiverIO (Jaegle et al., 2022) as score network, an efficient Transformer that allows for sub-quadratic compute, as well as FlashAttention (Dao et al., 2022) in implementation. Other efficient Transformer architectures and tricks like Jabri et al. (2022) can be used to improve training efficiency. The other factor is computational cost during inference. In MCF, we iterate 1000 timesteps to sample a conformer following DDPM (Ho et al., 2020). Experiments in Section 5.5 show that efficient sampling strategies, i.e.  DDIM (Song et al., 2021), can help significantly increase inference efficiency while maintain high-quality in sampled conformers. Other efficient variants of diffusion models like consistency model (Song et al., 2023) as well as distillation approaches (Berthelot et al., 2023) may be adapted to further decrease the sampling to single step. Also, recent works have demonstrated that diffusion generative model can generate samples following Boltzmann distributions when provided with Boltzmann-distributed training data (Arts et al., 2023). Driven by this, our proposed MCF can be adapted to generate molecular conformers that follow Boltzmann distributions when trained with corresponding data. Besides, recent flow matching generative model (Lipman et al., 2022) provides the flexibility of mapping between arbitrary distributions and access to exact log-likelihood estimation. Integrating flow matching framework could help sample molecular conformer from Boltzmann distribution instead of standard Gaussian. Some recent works (Flam-Shepherd & Aspuru-Guzik, 2023; O Pinheiro et al., 2024; Gruver et al., 2022) also show that expressive models can learn equivariance from data, but they have not thoroughly investigated molecular conformer generation.
Another limitation could be how well MCF performs in low data regime. The propsed method may not perform as well as conformer generation when applied to problems with limited data or related to sequential problems like molecular dynamics (MD) simulations. In future work, we plan to extend MCF to conditional inference. For example, molecular docking can be formulated as conformer generation problem conditioned on proteins (Corso et al., 2022). Also, current framework can be expanded to de novo drug designs where no molecule information is provided (Hoogeboom et al., 2022). Besides, scaling up our model to large molecules, like proteins, can be of great interest. MCF by nature provides the flexibility to generate from partially observed sample, which can be suitable for designing proteins with known functional motifs (Watson et al., 2023).
@@ A.2 Implementation details
In this section we describe implementation details for all our experiments. We also provide hyper-parameters and settings for the implementation of the score field network $\epsilon_{\theta}$ and compute used for each experiment in the paper. In our experiments, we split GEOM-QM9 and GEOM-DRUGS randomly based on molecules into train/validation/test (80%/10%/10%). At the end, for each dataset, we report the performance on 1000 test molecules. Thus, the splits contain 106586/13323/1000 and 243473/30433/1000 molecules for GEOM-QM9 and GEOM-DRUGS, respectively. We follow the exact same training splits for all baselines (Ganea et al., 2021; Jing et al., 2022).
@@ A.2.1 Score Field Network implementation details
The time-step $t$ is incorporated into the score computation by concatenating a positional embedding representation of $t$ to the context and query sets. The specific PerceiverIO settings used in all quantitatively evaluated experiments are presented in Tab. 4. An AdamW (Loshchilov & Hutter, 2017) optimizer is employed during training with a learning rate of $41e-$ 4. Cosine learning rate decay is deployed with 30K warmup steps. We use EMA with a decay of $0.999$. Models are trained for 300K steps on GEOM-QM9 and 750K steps on GEOM-DRUSG. All models use an effective batch size of 512. A modified version of the publicly available repository is used for PerceiverIO  [^1] https://huggingface.co/docs/transformers/model_doc/perceiver. Since molecules have different number of atoms, we set the number of context and query sets as the number of atoms during training and inference.
Hyper-parameter Small Base Large
num_freq_pos_embed $128$ $128$ $128$
num_latent $128$ $512$ $1024$
d_latennt $256$ $512$ $1024$
d_model $512$ $1024$ $1024$
num_enc_block $6$ $8$ $12$
num_dec_block $2$ $2$ $2$
num_self_attn_per_block $2$ $2$ $2$
num_self_attn_head $4$ $4$ $8$
num_cross_attn_head $4$ $4$ $8$
# param 13M 64M 242M
@@ A.2.3 Compute
For GEOM-QM9, we train models using a machine with 4 Nvidia A100 GPUs using precision BF16. For GEOM-DRUGS, we train models using precision FP32, where MCF-B is trained with 8 Nvidia A100 GPUs and MCF-L is trained with 16 Nvidia A100 GPUs.
@@ A.3.1 Ablation experiments
In this section we provide an ablation study over the key design choices of MCF. We run all our ablation experiments on the GEOM-QM9 dataset following the settings in GeoMol (Ganea et al., 2021) and Torsional Diff. (Jing et al., 2022). In particular we study: (i) how does performance behave as a function of the number of Laplacian eigenvectors used in $\varphi(v)$. (ii) How does the model perform without atom features (e.g. how predictable conformers are given only the graph topology, without using atom features). Results in Tab. 6 show that the graph topology $\mathcal{G}$ encodes a surprising amount of information for sampling reasonable conformers in GEOM-QM9, as shown in row 2. In addition, we show how performance of MCF changes as a function of the number of eigen-functions $k$. Interestingly, with as few as $k=2$ eigen-functions MCF is able to generate consistent accurate conformer.
@@ A.3.2 Architectural choices
To further investigate the design choices of architecture in proposed MCF, we include additional experiments on GEOM-QM9 as shown in Tab. 6. To investigate the effectiveness of using Laplacian eigenvectors from LBO eigen-decomposition as positional embedding, we leverage SignNet (Lim et al., 2022) as the positional embedding, which explicitly models symmetries in eigenvectors. Using SignNet does not benefit the performance when compared with the standard MCF. Though adding edge attributes in SignNet achieves better performance than SignNet alone, the performance is still not rival. Also, it’s worth mentioning that SignNet includes graph neural networks (Xu et al., 2018) and Set Transformer (Lee et al., 2019) which makes training less efficient.
In addition, we also report results using a vanilla Transformer encoder-decoder (TF) (Vaswani et al., 2017) as the backbone instead of PerceiverIO (PIO) (Jaegle et al., 2022). TF-base model contains 6 encoder layers and 6 decoder layers with 4 attention heads while TF-large contains 12 encoder layers and 12 decoder layers. The number of parameters match approximately with base and large sized PerceiverIO investigated in this work. Tab. 6 shows that TF-base is performing significantly worse than PIO-base with similar number of parameters. When increasing the model size, TF-large achieves on par performance as PIO-base, which validates the design choice of architecture in MCF.
@@
COV $\uparrow$ AMR $\downarrow$ COV $\uparrow$ AMR $\downarrow$
$k$ atom feat. PE backbone mean median mean median mean median mean median
28 YES LBO PIO-base 95.00 100.00 0.103 0.044 93.67 100.00 0.119 0.055
28 NO LBO PIO-base 90.70 100.00 0.187 0.124 79.82 93.86 0.295 0.213
16 YES LBO PIO-base 94.87 100.00 0.139 0.093 87.54 100.00 0.220 0.151
8 YES LBO PIO-base 94.28 100.00 0.162 0.109 84.27 100.00 0.261 0.208
4 YES LBO PIO-base 94.57 100.00 0.145 0.093 86.83 100.00 0.225 0.151
2 YES LBO PIO-base 93.15 100.00 0.152 0.088 86.97 100.00 0.211 0.138
28 YES SignNet PIO-base 94.10 100.00 0.153 0.098 87.50 100.0 0.222 0.152
28 YES SignNetattr PIO-base 95.30 100.00 0.143 0.091 90.20 100.00 0.197 0.135
28 YES LBO TF-base 94.92 100.00 0.131 0.083 89.33 100.00 0.194 0.132
28 YES LBO TF-large 95.49 100.00 0.110 0.061 93.48 100.00 0.135 0.073
Table 6: Ablation study with different network architectures on GEOM-QM9.
@@ A.3.3 Ensemble properties
To fully assess the quality of generated conformers we also compute chemical property resemblance between the synthesized and the authentic ground truth ensembles. We select a random group of 100 molecules from the GEOM-DRUGS and produce a minimum of $2K$ and a maximum of $32$ conformers for each molecule following (Jing et al., 2022). Subsequently, we undertake a comparison of the Boltzmann-weighted attributes of the created and the true ensembles. To elaborate, we calculate the following characteristics using xTB (as documented by (Bannwarth et al., 2019)): energy ($E$), dipole moment ($\mu$), the gap between HOMO and LUMO ($\Delta\epsilon$), and the lowest possible energy, denoted as $E_{\text{min}}$. Since we don’t have the access to the exact subset of DRUGS used in (Jing et al., 2022), we randomly pick three subsets and report the averaged and standard deviation over three individual runs with different random seeds. The results are listed in Tab. 7. Our model achieves the lowest error on $E_{\text{min}}$ when compared with other baselines, which demonstrates that MCF succeeds at generating stable conformers that are very close to the ground states. This could root from the fact that MCF doesn’t rely on rule-based cheminfomatics methods and the model learns to better model stable conformers from data. Besides, MCF achieves competitive performance on $\mu$ and $\Delta\epsilon$. However, the error of $E$ is high compared to the rest of approaches, meaning that though MCF performs well in generating samples close to ground states, it may also generate conformers with high energy that are not plausible in the dataset.
@@
OMEGA 0.68 0.66 0.68 0.69
GeoDiff 0.31 0.35 0.89 0.39
GeoMol 0.42 0.34 0.59 0.40
Tor. Diff. 0.22 0.35 0.54 0.13
MCF 0.68$\pm$0.06 0.28$\pm$0.05 0.63$\pm$0.05 0.04$\pm$0.00
Tor. Diff. (our eval) 3.07$\pm$2.32 0.61$\pm$0.38 1.71$\pm$1.69 4.11$\pm$7.91
MCF 1.00$\pm$0.70 0.44$\pm$0.36 1.32$\pm$1.40 1.16$\pm$2.02
@@
Figure 7: Examples of conformers with lowest and highest energies in ground truth, MCF samples, and Torsional Diff. samples for different molecules.
@@
To further evaluate the performance on ensemble properties, we randomly pick 10 molecules from test set of GEOM-DRUGS and compare MCF with our replica Torsional Diff. on the subset as shown in the last two rows of Tab. 7. We use the checkpoints from the public GitHub repository [^2] https://github.com/gcorso/torsional-diffusion of Torsional Diff. to sample conformers. Unlike previous setting which only sample 32 conformers, we sample $2K$ conformers for a molecule with $K$ ground truth conformers. We report the average and standard deviation of errors over the 10 molecules. It is indicated that MCF generates samples with ensemble properties that are closer to the ground truth. Fig. 7 shows the conformers with lowest and highest energy in ground truth, MCF samples, and Torsional Diff. samples.
A.4 Continuous conformers
Molecular conformers are defined as discrete atomic positions in 3D Euclidean space. Since MCF encodes continuous conformer field, it can be continuously evaluated in $\mathcal{G}$, which maps arbitrary points in $\mathcal{G}$ to 3D positional space $\mathbbm{R}^{3}$. In order to do this, for a point $p$ in a bond connecting atoms $(v_{i},v_{j})$ we linearly interpolate the Laplacian eigenvector representation of it’s endpoints $\varphi(p)=\alpha\varphi(v_{i})+(1-\alpha)\varphi({v_{j}})$, we then feed this interpolated Laplacian eigenvector into the model to sample its 3D position in the conformer field. We visualize results in Fig. 1 and 8. We generated this visualizations an MCF model trained on GEOM-QM9 without atom features. Note that while MCF is never trained on points along molecular bonds, it manages to generate plausible 3D positions for such points.
A.5 Additional visualization
Fig. 9 show some examples of sampled conformers from MCF with different sampling steps. It is illustrated that even with very limited sample steps, MCF can still generate plausible conformers especially for the heavy atoms. We also show examples of conformers from ground truth, Torsional Diff., and our MCF. Fig. 10 and 11 depict samples from GEOM-DRUGS and GEOM-XL, respectively. We found the samples that are most aligned with ground truth and plot them side by side.